# Cross-Service Threat Intelligence in LLM Services using Privacy-Preserving Fingerprints

### v1

@@ 3.2.1 Empirical Analysis of Randomized Response Impact
We quantify the impact of randomized response mechanism (Warner, 1965) noise on detection utility by sweeping the per-bit privacy budget $\alpha\!\in\![0.25,3.0]$ (Algorithm 1, Step 4) on the most challenging paraphrase attack variant. For each $\alpha$, a bit in the 768-dimensional binary embedding is kept with probability $p=\tfrac{e^{\alpha}}{e^{\alpha}+1}$ and flipped otherwise, yielding an expected self-Hamming distortion (1p)d(1-p)d, where $d$ is the fingerprint dimension (here, $d=768$). At $\alpha=0.25$, $p=0.562$ and the mechanism flips $\approx 0.438\times 768\approx 336$ bits (i.e., the whole binary vector is approximately completely randomized). As shown in the top-right confusion matrix of Figure 8, at this extreme privacy setting there are 821 false positives and 97 false negatives, resulting in an F1-score of only 0.65. Such high noise levels provide maximal privacy but render the binary fingerprints nearly random (Figure 9), offering little practical utility for threat detection.
@@
Randomized Response and Hamming Distance. To empirically validate BinaryShield’s privacy model, we conducted a controlled experiment measuring the actual Hamming distance introduced by the BinaryShield’s randomized response mechanism.
To this end, for 500 prompts from our evaluation dataset, BinaryShield generated 768-dimensional binary fingerprints and applied its differential privacy transformation by varying $\alpha$ values from 0.2 to 3.4. For each configuration, we computed the Hamming distance between original and noise-perturbed fingerprints, comparing observed distances against the theoretical expectation (1p)d(1-p)d. As a baseline reference, we generated 1,000 pairs of independent random binary vectors to establish the average distance for completely uncorrelated binary vectors. The mean Hamming distance for these independent vectors is $384.44\pm 13.65$, which serves as a baseline for comparison. Figure 9 demonstrates that our theoretical model precisely predicts the observed self-Hamming distortion across all privacy levels confirming BinaryShield’s alignment with theory.
@@
At $\alpha=0.2$, BinaryShield’s perturbed binary fingerprint reaches a Hamming distance of $346\pm 13.41$, nearly matching the random baseline. This results in lower F1 scores at small $\alpha$ (Figure 8).
@@
Figure 10: Scalability of privacy-preserving threat correlation. Accuracy@k for BinaryShield vs non-private dense embeddings baseline across hybrid corpora (10K-100K). Accuracy remains flat with scale; BinaryShield retains 93% of baseline Accuracy@1 while providing privacy.
3.3 Real-World Deployment Analysis
While earlier sections validate BinaryShield supporior performance as compared to privacy preserving baseline (SimHash), we now turn to evaluating BinaryShield’s effectiveness in realistic cross-service threat intelligence scenarios. Specifically, it must be shown that privacy-preserving correlation remains effective when (i) malicious prompts are sparse relative to benign traffic and (ii) corpus size scales by an order of magnitude (iii) Privacy-utility trade-off is maintained, i.e., the accuracy gap between BinaryShield and non-private dense embeddings remains bounded, and (iv) efficiency is maintained in terms of computational and storage overhead.
We study this setting using hybrid corpora that interleave a (fixed) malicious prompt injection (V-Paraphrase) with large volumes of real user interactions from WildChat (Zhao et al., 2024), thereby emulating an enterprise log in which only a small fraction of entries are prompt injection attacks.
3.3.1 Large-Scale Threat Intelligence Scalability
The goal is to quantify how BinaryShield’s privacy-preserving fingerprint correlation accuracy behaves as the total corpus grows from 10K to 100K entries, and characterize the accuracy gap relative to a non-private dense embedding baseline. In this experiment, BinaryShield uses OpenAI’s text-embedding-3-large to generate dense embeddings.
Figure 11: Accuracy@k gap (Dense Embeddings - BinaryShield) across corpus sizes. Rapid narrowing for higher k demonstrate that binary quantization plus local DP introduce a constant, not compounding, utility cost.
For each target corpus size $C\in\{10\text{K},20\text{K},\dots,100\text{K}\}$, we construct a hybrid dataset by injecting a fixed malicious prompt set into CMC-M benign WildChat prompts (where $M$ is the number of distinct attack prompts). Each malicious prompt is issued as a query; success is recorded if its exact counterpart (or semantically identical variant, depending on $k$) appears within the top-$k$ candidates when ranked by Hamming distance over BinaryShield’s fingerprints, or cosine similarity for the dense baseline. We fix the local differential privacy parameter at $\alpha=2$ (different $\alpha$ values are also evaluated in Section 3.3.2).
@@ 3.3.1 Large-Scale Threat Intelligence Scalability
Furthermore, Figure 11 shows that the Accuracy@k gaps between BinaryShield and the dense baseline with X-axis showing corpus sizes (10K to 100K) and Y-axis showing the gap at different top-k accuracy. The gap remains tightly bounded: it ranges from 5.31 to 6.15 percentage points (mean 5.66% points) at $k=1$, indicating that BinaryShield retains over 93% of the dense baseline’s accuracy even under tight local differential privacy constraints. Crucially, the gap does not significantly widen with corpus size. Furthermore, increasing $k$ rapidly closes the gap, as shown in Figure 10 and Figure 11. For instance at 100K entries, BinaryShield achieves 79.2% Accuracy@1, 90.8% Accuracy@3, and 92.9% Accuracy@5. On the same 100K corpus, the dense embeddings achieve 85.3% Accuracy@1, 94.7% Accuracy@3, and 96.5% Accuracy@5. The gap at $k=5$ is only 3.6% points, indicating that BinaryShield’s with privacy feature reaches close to the non-private baseline of dense embeddings. This confirms that the semantic signal preserved after (i) sign-only quantization and (ii) per-bit randomized response still remains sufficiently rich for practical threat correlation. For every corpus size, BinaryShield’s incremental gains from $k=1$ to $k=2$ average roughly +8 percentage points (Figure 10). This reflects that most residual misses at $k=1$ are near-boundary semantic neighbors retrievable with minimal expansion. Subsequent marginal gains ($k>2$) diminish smoothly. Figure 11 shows the gap curves monotonically compress as $k$ increases. The gap is attributable to two irreversible transformations: (a) many-to-one sign mapping (loss of magnitude information) and (b) randomized response bit flips calibrated by $\alpha=2$. That the gap does not drift upward with scale empirically supports that these transformations in BinaryShield behave as scale-neutral perturbations rather than compounding sources of semantic erosion. In an operational setting where an analyst (or automated response engine) can examine the top 3–5 correlated candidates, BinaryShield delivers privacy-preserving threat intelligence with accuracy comparable to a fully non-private dense embedding pipeline. This constitutes a strong utility guarantee under stringent local privacy constraints and validates BinaryShield’s suitability for multi-service deployment where corpus sizes routinely exceed tens of thousands of daily log entries.
@@
Summary. BinaryShield achieves over 93% accuracy of dense embeddings (non-private baseline) and remains scalable in practical deployment settings with randomized response mechanism, demonstrating robust privacy-preserving threat correlation at enterprise scale.
3.3.2 BinaryShield’s Privacy Parameter Performance at Scale
@@
Figure 12: Privacy-utility trade-off for BinaryShield across $\alpha$ values on 50K hybrid corpus. Accuracy@1 (blue line) and Accuracy@5 (dotted black line) show phase transition behavior as privacy budget increases. The non-private dense embedding baseline is shown as horizontal green and orange lines at 86.04% (Accuracy@1) and 96.56% (Accuracy@5) respectively.
@@
The relationship between privacy protection and threat detection utility at scale represents a fundamental tension in the cross-boundary threat intelligence system. Similar to Section 3.2.1, to characterize this trade-off precisely, we evaluate BinaryShield’s performance across a spectrum of privacy budgets ($\alpha$) on a fixed 50K-entry hybrid corpus containing both malicious prompts and benign WildChat interactions (Section 3.3.1). The privacy parameter $\alpha$ directly controls the randomized response mechanism through the bit preservation probability $p=e^{\alpha}/(e^{\alpha}+1)$, with the corresponding bit flip probability being 1p1-p. Smaller $\alpha$ values provide stronger privacy guarantees by increasing the probability of bit flips in the binary fingerprint, while larger values preserve more semantic information at the cost of reduced privacy protection.
@@
The results show a smooth transition in the privacy-utility trade-off, especially between $\alpha=0.75$ and $\alpha=2.0$. For significantly high-privacy needs, $\alpha=1.5$ gives strong privacy. Since organizations have already strict security policies, for internal sharing, relaxing privacy to $\alpha=2.0$$2.5$ yields near-optimal utility with realistic meaningful privacy protection. Overall, BinaryShield potentially can operate in a wide range of privacy-utility trade-offs.
@@
Summary. BinaryShield empowers organizations to balance privacy and utility as needed, consistently delivering strong accuracy and realistic privacy protection. The organizations can flexibly tune the privacy-utility trade-off based on their operational needs, regulatory requirements, and threat detection scenarios.
3.3.3 BinaryShield Computational and Storage Efficiency Analysis
@@
The operational viability of any threat intelligence system fundamentally depends on its ability to process vast quantities of data with minimal latency and storage overhead. Historical lessons from traditional cybersecurity infrastructure underscore this requirement. One of the major reason that signature-based ecosystems (antivirus, IDS) scaled is because compact hashes (e.g., MD5, fuzzy-hash families). Similarly, modern LLM services face an even more daunting challenge, processing billions of queries daily across distributed infrastructure. For BinaryShield to enable practical cross-boundary threat intelligence, it must achieve orders-of-magnitude improvements in both computational speed and storage efficiency compared to non-private dense embeddings. Every extra byte per fingerprint multiplies into material storage cost, and every millisecond of per-fingerprint scan latency stretches incident response time.
@@
Figure 13(a) reveals the stark computational advantages of BinaryShield’s binary fingerprints over dense embeddings. At a corpus size of 10K entries, dense embeddings require 0.87 seconds for similarity search of 968 prompt injections, while BinaryShield completes the same operation in just 0.032 seconds, a 26.7x speedup. This performance gap widens dramatically as corpus size increases. At 100K entries, dense embeddings demand 14.52 seconds, rendering real-time threat correlation infeasible for high-volume services. In contrast, BinaryShield maintains sub-second performance at 0.38 seconds, achieving a 38.1x speedup. The computational efficiency stems from the fundamental difference in similarity computation. Dense embeddings require floating-point dot products or cosine similarity calculations across thousands of dimensions, operations that scale poorly even with optimized linear algebra libraries. BinaryShield’s binary fingerprints enable Hamming distance computation through simple XOR operations.
@@
(a) Search time comparison between non-private dense embeddings and BinaryShield across corpus sizes. Note the logarithmic scale on the Y-axis. Dense embeddings exhibit super-linear growth reaching 16.3 seconds at 80K entries, while BinaryShield maintains sub-second performance throughout, demonstrating its suitability for real-time threat correlation at enterprise scale.
(b) Storage size comparison demonstrating BinaryShield’s 64x reduction compared to dense embeddings. Dense embeddings scale from 256.88 MB at 10K entries to 2,366.25 MB at 100K entries, while BinaryShield requires only 4.01 MB to 36.97 MB respectively, enabling cost-effective long-term threat intelligence retention.
Figure 13: Computational and storage efficiency analysis of BinaryShield versus dense embeddings.
The storage requirements present an equally compelling case for BinaryShield’s architecture, as illustrated in Figure 13(b). Dense embeddings consume 256.88 MB for 10K entries, growing linearly to 2,366.25 MB at 100K entries. This represents a substantial infrastructure burden when organizations must maintain threat intelligence databases spanning billions of historical queries across multiple compliance boundaries. BinaryShield reduces storage requirements by a factor of 64, requiring only 4.01 MB for 10K entries and 36.97 MB for 100K entries. This 64x reduction directly translates from the binary quantization process, where each embedding dimension is compressed from a floating-point value to a single bit, followed by the addition of noise through randomized response that preserves the binary nature.
@@
The practical implications of these efficiency gains extend beyond raw performance metrics (Section  3.2 and  3.3.1). Consider an organization processing 100 million queries daily across ten compliance boundaries. With dense embeddings, maintaining a rolling 30-day threat intelligence window would require approximately 7.1 TB of storage per boundary and dedicated GPU clusters for similarity search making it near to impractical choice. BinaryShield reduces this to 111 GB per boundary, fitting comfortably in memory on commodity servers while enabling CPU-based similarity search at wire speed. This efficiency democratizes threat intelligence capabilities, allowing even resource-constrained services to participate in collaborative defense without significant infrastructure investment.
@@
Furthermore, the computational efficiency enables new operational capabilities previously infeasible with dense embeddings. Security teams can now perform retrospective threat hunting across months of historical data in minutes rather than hours. The reduced storage footprint also facilitates comprehensive threat intelligence archival, enabling long-term trend analysis and attribution of persistent attack campaigns that evolve over extended periods.
@@
Summary. BinaryShield’s computational and storage efficiency fundamentally transforms cross-boundary threat intelligence economics. By reducing search latency up to 38x and storage requirements up to 64x while maintaining 93% of non-private baseline top-1 accuracy with privacy, BinaryShield enables enterprise-scale threat correlation. This positions BinaryShield as a practical foundation for collaborative threat intelligence systems against prompt injection attacks that respect regulatory boundaries while enabling effective cross-service intelligence sharing.
4 Related Work and Discussion
Prompt injection attacks pose significant risks to LLMs by embedding covert instructions within benign-looking input data to manipulate model outputs. A number of recent studies have explored both the attack methods and potential defenses against these vulnerabilities.
4.1 Attacks
[…]
4.2 Attack Defenses
Recent work proposes a variety of defense strategies against prompt injection attacks. These approaches differ in their methods and operating settings, ranging from inference-time countermeasures to white-box techniques and architectural safeguards.

### v2

@@ III-B1 Empirical Analysis of Randomized Response Impact
We quantify the impact of randomized response mechanism [51] noise on detection utility by sweeping the per‑bit privacy budget $\alpha\!\in\![0.25,3.0]$ (Algorithm 1, Step 4) on the most challenging paraphrase attack variant. For each $\alpha$, a bit in the 768‑dimensional binary embedding is kept with probability $p=\tfrac{e^{\alpha}}{e^{\alpha}+1}$ and flipped otherwise, yielding an expected self‑Hamming distortion (1p)d(1-p)d, where $d$ is the fingerprint dimension (here, $d=768$). At $\alpha=0.25$, $p=0.562$ and the mechanism flips $\approx 0.438\times 768\approx 336$ bits (i.e., the whole binary vector is approximately completely randomized). As shown in the top-right confusion matrix of Figure 8, at this extreme privacy setting there are 821 false positives and 97 false negatives, resulting in an F1-score of only 0.65. Such high noise levels provide maximal privacy but render the binary fingerprints nearly random (Figure 9), offering little practical utility for threat detection.
@@
Randomized Response and Hamming Distance. To empirically validate BinaryShield’s privacy model, we conducted a controlled experiment measuring the actual Hamming distance introduced by the BinaryShield’s differential privacy.
To this end, for 500 prompts from our evaluation dataset, BinaryShield generated 768-dimensional binary fingerprints and applied its differential privacy transformation by varying $\alpha$ values from 0.2 to 3.4. For each configuration, we computed the Hamming distance between original and noise-perturbed fingerprints, comparing observed distances against the theoretical expectation (1p)d(1-p)d. As a baseline reference, we generated 1,000 pairs of independent random binary vectors to establish the average distance for completely uncorrelated binary vectors. The mean Hamming distance for these independent vectors is $384.44\pm 13.65$, which serves as a baseline for comparison. Figure 9 demonstrates that our theoretical model precisely predicts the observed self-Hamming distortion across all privacy levels confirming BinaryShield’s alignment with theory. At $\alpha=0.2$, BinaryShield’s perturbed binary fingerprint reaches a Hamming distance of $346\pm 13.41$, nearly matching the random baseline. This results in lower F1 scores at small $\alpha$ (Figure 8).
@@
Figure 10: Scalability of BinaryShield. Accuracy@k for BinaryShield vs non-private dense embeddings baseline across hybrid corpora (10K-100K). BinaryShield retains 93% of baseline Accuracy@1 while providing privacy.
III-C Real-World Deployment Analysis
While earlier sections validate BinaryShield’s superior performance as compared to the privacy-preserving baseline (SimHash), we now turn to evaluating BinaryShield’s effectiveness in realistic cross-service threat intelligence scenarios. Specifically, it must be shown that privacy-preserving correlation remains effective when (i) malicious prompts are sparse relative to benign traffic and (ii) corpus size scales by an order of magnitude (iii) privacy-utility trade-off is maintained, i.e., the accuracy gap between BinaryShield and non-private dense embeddings remains bounded, and (iv) efficiency is maintained in terms of computational and storage overhead.
III-C1 Large-Scale Threat Intelligence Scalability
To simulate realistic enterprise conditions, we query attack fingerprints against corpora where benign prompts vastly outnumber malicious ones, measuring Accuracy@k as successful detection when the true variant ranks within the top-k retrieved candidates. The goal is to quantify how BinaryShield’s privacy-preserving fingerprint correlation accuracy behaves as the total corpus grows, and characterize the accuracy gap relative to a non-private dense embedding baseline. In this experiment, BinaryShield uses OpenAI’s text-embedding-3-large to generate embeddings.
We study this setting using hybrid corpora that interleave a (fixed) malicious prompt injection (V-Paraphrase) with large volumes of real user interactions from WildChat [58], thereby emulating an enterprise log in which only a small fraction of entries are attacks. For each target corpus size $C\in\{10\text{K},20\text{K},\dots,100\text{K}\}$, we construct a hybrid dataset by injecting a fixed malicious prompt set into CMC-M benign WildChat prompts (where $M$ is the number of distinct attack prompts). Each malicious prompt is issued as a query; success is recorded if its exact counterpart (or semantically identical variant, depending on $k$) appears within the top-$k$ candidates when ranked by Hamming distance over BinaryShield’s fingerprints, or cosine similarity for the dense baseline. We fix the local differential privacy parameter at $\alpha=2$ (different $\alpha$ values are also evaluated in Section III-C2).
@@ III-C1 Large-Scale Threat Intelligence Scalability
Furthermore, Figure˜13 (in the appendix) shows that the Accuracy@k gaps between BinaryShield and the dense baseline with X-axis showing corpus sizes (10K to 100K) and Y-axis showing the gap at different top-k accuracy. The gap remains tightly bounded: it ranges from 5.31% to 6.15% (mean 5.66% points) at $k=1$, indicating that BinaryShield retains over 93% of the dense baseline’s accuracy even under tight local differential privacy constraints. Crucially, the gap does not significantly widen with corpus size. Furthermore, increasing $k$ rapidly closes the gap, as shown in Figure 10 and Figure 13. For instance at 100K entries, BinaryShield achieves 79.2% Accuracy@1, 90.8% Accuracy@3, and 92.9% Accuracy@5. On the same 100K corpus, the dense embeddings achieve 85.3% Accuracy@1, 94.7% Accuracy@3, and 96.5% Accuracy@5. The gap at $k=5$ is only 3.6% points, indicating that BinaryShield’s with privacy feature reaches close to the non-private baseline of dense embeddings. This confirms that the semantic signal preserved after (i) sign-only quantization and (ii) per-bit randomized response still remains sufficiently rich for practical threat correlation. For every corpus size, BinaryShield’s incremental gains from $k=1$ to $k=2$ average roughly +8% (Figure 10). This reflects that most residual misses at $k=1$ are near-boundary semantic neighbors retrievable with minimal expansion. Subsequent marginal gains ($k>2$) diminish smoothly. Figure 13 shows the gap curves monotonically compress as $k$ increases. The gap is attributable to two irreversible transformations: (a) many-to-one sign mapping (loss of magnitude information) and (b) randomized response bit flips calibrated by $\alpha=2$. That the gap does not drift upward with scale empirically supports that these transformations in BinaryShield behave as scale-neutral perturbations rather than compounding sources of semantic erosion. In an operational setting where an analyst (or automated response engine) can examine the top 3–5 correlated candidates, BinaryShield delivers privacy-preserving threat intelligence with accuracy comparable to a fully non-private dense embedding pipeline. This constitutes strong utility under privacy constraints and validates BinaryShield’s suitability for multi-service deployment.
@@
Summary. BinaryShield retains over 93% accuracy of dense embeddings (non-private baseline) and remains scalable in practical deployment settings with randomized response mechanism, demonstrating robust privacy-preserving threat correlation at enterprise scale.
III-C2 BinaryShield’s Privacy Parameter Performance at Scale
@@
Figure 11: Privacy-utility trade-off for BinaryShield across $\alpha$ values on 50K hybrid corpus. Accuracy@1 (blue line) and Accuracy@5 (black line) show phase transition behavior as privacy budget increases.
@@
The relationship between privacy protection and threat detection utility at scale represents a fundamental tension in the cross-boundary threat intelligence system. Similar to Section III-B1, to characterize this trade-off precisely, we evaluate BinaryShield’s performance across a spectrum of privacy budgets ($\alpha$) on a fixed 50K-entry hybrid corpus containing both malicious prompts and benign WildChat interactions (Section III-C1). The privacy parameter $\alpha$ directly controls the randomized response mechanism through the bit preservation probability $p=e^{\alpha}/(e^{\alpha}+1)$, with the corresponding bit flip probability being 1p1-p. Smaller $\alpha$ values provide stronger privacy guarantees by increasing the probability of bit flips in the binary fingerprint, while larger values preserve more semantic information at the cost of reduced privacy protection.
@@
The results show a smooth transition in the privacy-utility trade-off, especially between $\alpha=0.75$ and $\alpha=2.0$. Since organizations have already strict security policies, for internal sharing, relaxing privacy to $\alpha=2.0$$2.5$ yields near-optimal utility with realistic meaningful privacy protection. Overall, BinaryShield potentially can operate in a wide range of privacy-utility trade-offs.
@@
Summary. BinaryShield empowers organizations to balance privacy and utility as needed, consistently delivering strong accuracy and realistic privacy protection. The organizations can flexibly tune the privacy-utility trade-off based on their operational and regulatory needs.
III-C3 BinaryShield Computational Efficiency Analysis
@@
The operational viability of any threat intelligence system fundamentally depends on its ability to process vast quantities of data with minimal latency and storage overhead. Historical lessons from traditional cybersecurity infrastructure underscore this requirement. One of the major reasons that signature-based ecosystems (antivirus, IDS) scaled is because compact hashes (e.g., MD5, fuzzy-hash families). Similarly, modern LLM services face an even more daunting challenge, processing billions of queries daily across distributed infrastructure. For BinaryShield to enable practical cross-boundary threat intelligence, it must achieve orders-of-magnitude improvements in both computational speed and storage efficiency compared to non-private dense embeddings. Every extra byte per fingerprint multiplies into material storage cost, and every millisecond of per-fingerprint scan latency stretches incident response time.
@@
Figure 12 reveals computational advantages of BinaryShield’s fingerprints over dense embeddings. At a corpus size of 10K entries, dense embeddings require 0.87 seconds for similarity search of 968 prompt injections, while BinaryShield completes the same operation in just 0.032 seconds, a 26.7x speedup. This performance gap widens dramatically as corpus size increases. At 100K entries, dense embeddings demand 14.52 seconds, rendering real-time threat correlation infeasible for high-volume services. In contrast, BinaryShield maintains sub-second performance at 0.38 seconds, achieving a 38.1x speedup. The computational efficiency stems from the fundamental difference in similarity computation. Dense embeddings require floating-point dot products or cosine similarity calculations, operations that scale poorly even with optimized linear algebra libraries. BinaryShield’s binary fingerprints enable Hamming distance computation through simple XOR operations. We discuss BinaryShield’s storage utilization in the appendix.
@@
Figure 12: Search time comparison between non-private dense embeddings and BinaryShield across corpus sizes. Note the logarithmic scale on the Y-axis. BinaryShield maintains sub-second performance, demonstrating its suitability for real-time threat correlation at enterprise scale.
@@
The practical implications of these efficiency gains extend beyond raw performance metrics (Section  III-B and  III-C1). Consider an organization processing 100 million queries daily across ten services. With dense embeddings, maintaining a rolling 30-day threat intelligence window would require approximately 7.1 TB of storage per boundary and dedicated GPU clusters for similarity search making it near to impractical choice. BinaryShield reduces this to 111 GB per boundary, fitting comfortably in memory on commodity servers while enabling CPU-based similarity search. This efficiency democratizes threat intelligence capabilities, allowing even resource-constrained services to participate in collaborative defense without significant infrastructure investment.
@@
Furthermore, the computational efficiency enables new operational capabilities. Security teams can now perform retrospective threat hunting across months of historical data in minutes rather than hours. The reduced storage footprint also facilitates comprehensive threat intelligence archival, enabling long-term trend analysis and attribution of persistent attack campaigns that evolve over extended periods.
@@
Summary. BinaryShield’s computational and storage efficiency fundamentally transforms cross-boundary threat intelligence economics. By reducing search latency up to 38x and storage requirements while maintaining 93% of non-private baseline top-1 accuracy with privacy, BinaryShield enables enterprise-scale threat correlation. This positions BinaryShield as a practical foundation for collaborative threat intelligence systems against prompt injection attacks.
IV Related Work and Discussion
Attacks. Several works have focused on systematically characterizing prompt injection attacks. For instance, [24] demonstrate how hidden instructions can be inserted into external content to trigger harmful model behaviors. Other works [52, 60, 57, 14, 28, 33, 45, 15, 11] further contribute to our understanding by exploring a variety of prompt injection scenarios. In addition, [54] propose SOS, a training-time “soft prompt” attack that implants backdoors into open-source LLMs. The LLM behaves normally, until a trigger token activates malicious behaviors like jailbreaks, prompt stealing, or output manipulation. In contrast, [44] introduce Crescendo, an inference-time jailbreak attack that unfolds over multiple dialogue turns. Supporting this observation, [9] performs an in-depth analysis of such multi-turn jailbreak attacks and show why defenses designed for single-turn interactions, like circuit breakers [59], may be insufficient. By examining internal model representations, [9] shows that multi-turn prompts can gradually reframe harmful outputs as safe, effectively evading safeguards designed for single-turn interactions. BIPIA [55] systematically assesses the risks posed by malicious instructions embedded in third-party content and introduces two defense paradigms: one based on prompt engineering and in-context strategies in a black-box setting, and another using adversarial training with special tokens in a white-box setting. Finally, [2] introduces a realistic testbed for indirect prompt injection attacks targeting an LLM-based email assistant. Collectively, these studies [24, 52, 60, 57, 14, 28, 33, 45, 15, 11, 54, 44, 9, 55, 2] provide a comprehensive overview of the diverse strategies employed to exploit prompt injection attacks, underscoring the ongoing need for robust security measures in LLM-integrated applications and services.